# LLM-Resistant Math Word Problem Generation via Adversarial Attacks

### v2

@@ Abstract
Large language models (LLMs) have significantly transformed the educational landscape. As current plagiarism detection tools struggle to keep pace with LLMs’ rapid advancements, the educational community faces the challenge of assessing students’ true problem-solving abilities in the presence of LLMs. In this work, we explore a new paradigm for ensuring fair evaluation—generating adversarial examples which preserve the structure and difficulty of the original questions aimed for assessment, but are unsolvable by LLMs. Focusing on the domain of math word problems, we leverage abstract syntax trees to structurally generate adversarial examples that cause LLMs to produce incorrect answers by simply editing the numeric values in the problems. We conduct experiments on various open- and closed-source LLMs, quantitatively and qualitatively demonstrating that our method significantly degrades their math problem-solving ability. We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models. Additionally, we conduct automatic analysis on math problems and investigate the cause of failure, offering a nuanced view into model’s limitation.
@@ 1 Introduction
Figure 1: Method Overview Given a MWP that an LLM can correctly solve, our method first transforms it into Python code. The Python code then is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem. We place constraints on the nodes of the AST to ensure that the modified problem maintains the same difficulty level as the original problem. Despite this, we find that the resulting adversarial examples cause LLMs to predict incorrect answers.
@@
In this work, we evaluate several major LLMs and demonstrate the effectiveness of our method by achieving a significant attack success rate (ASR). Our approach outperforms the previous rephrasing attack by an average of 60% ASR. We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models (e.g., GPT-4) by reducing API request calls by 90% without sacrificing performance. We conduct human evaluations to verify that our generated problems indeed preserve the original coherence and difficulty. Furthermore, we perform a regression analysis and find that our adversarial examples exploit different weaknesses of each model, offering valuable insights and a nuanced view into model’s limitation.
@@ 2 Background and Related Work
Our work also relates closely to LLMs math reasoning ability (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023). Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023a), such as chain-of-thought (CoT) (Wei et al., 2022). Converting the MWPs’ solving steps into code or symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023). [^1] While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve. Therefore, we try to minimize prompt engineering by using a unified prompt with zero-shot CoT for all evaluations. Prompt can be found in Appendix A.2.
@@
Adversarial attacks on LLMs involve modifying prompts to alter their behavior (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023). Similarly, we modify MWPs to cause LLMs to output incorrect answers. Bubeck et al. (2023) conducted limited memorization tests on MWPs by randomly changing numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not solely rely on memorization but apply general solution methods. However, their study’s sample size was relatively small, and we demonstrate that modifying numbers in MWPs causes LLMs to fail on a larger scale. On the other hand, Zhou et al. (2023) rephrased MWPs by changing words while preserving numeric values. This approach risks altering the original context and introducing inconsistencies and problematic content (see Appendix B), necessitating extensive human validation. Therefore, we focus on changing numerical values, preserving the underlying logic. To the best of our knowledge, this is the first study to undertake such an endeavor.
@@ 3 Methodology
MWPs are presented using natural language, which creates challenges for systematic structural and syntactic modifications. In this section, we describe our approach to generating adversarial MWPs. An overview of our method can be found in Figure 1. We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer. We define $G$ as a ground-truth function that computes the correct answer for a math problem [^2] In our case $G$ is generated Python code. and $F$ as a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise. We denote the set of all possible adversarial modifications to it as:
@@ 3.1 Mapping From MWPs to Code to Tree
To structurally modify MWPs, we first use GPT-4 to generate the Python code that reflects the solution steps given the problem and its final answer. [^3] We only require the problem and its final answer, which is almost always given in the context of math problems. Prompt can be found in Appendix A.1. As a sanity check, we select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015) and check if the generated program results in correct answers. Given the execution steps in code are almost always deterministic, encouragingly, we found only 1% and 2.5% generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively. We further investigate the generated Python code in §5.
Next, we utilize AST, a tree for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree representation. We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement. The final print statement that outputs the answer is the root of the tree. Each AST has mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes. We discuss the nodes in AST in detail in Appendix C.
@@ 3.2 Adversarial Example Generation
In an adversarial setting, attackers can perform multiple attacks on the system, and any successful attack signifies the possible flaws of the system (Carlini & Wagner, 2017). Similarly, we generate adversarial examples which are different versions of the original MWPs to attack LLMs.
@@
It’s essential to maintain the original difficulty and coherence of the MWPs despite changes in their numeric values. For instance, changing one-digit multiplication to four-digit multiplication significantly alters the problem’s complexity. To maintain mathematical logic intact, the order of magnitude of numbers in the problem should also remain unchanged. For example, changing a problem from “Jack ate 2 out of a total of 5 apples” to “Jack ate 5 out of a total of 2 apples” not only introduces the concept of negative numbers but also alters the problem’s logical structure and coherency. Furthermore, such modifications can lead to unrealistic scenarios that deviate from the intended learning objectives and may cause confusion.
@@
To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node to control the generation quality. Given the original node value $h$ and its new value $h^{\prime}$, a newly generated problem is valid if and only if $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes. A list of node constraints is:
@@
The node constraints ensure the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated. Different values for the variable nodes can also lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$.
To mitigate this, we propose three generation methods to distinguish generated adversarial examples with different levels of difficulty. We decompose the values $h\approx a\times 10^{b}$, where $a$ is an integer that is not divisible by 10 and with at most $c$ digits. Depending on the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
  • $M1$: Free Generation: Allowing the generated problem to have a wide range of numbers with minimal constraints, with each $h^{\prime}$ to be $a^{\prime}\times 10^{b}$. Regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$. If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$.
  • $M2$: Count of Digits: Constraining the generated value $h^{\prime}$ has the same number of digits as $h$. For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264. This ensures $h^{\prime}$ is always within a reasonable range relative to $h$ and maintains similar problem difficulty levels. For more variability, $h^{\prime}$ can range from 1 to 99 if $h$ has only one digit. For decimal variable values like $h=1.25$, we use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
  • $M3$: Count of Scientific Numbers: Constraining the new value shares a similar scientific digit count and is within a similar range to the original value. For example, for $h=1500$, $h^{\prime}\sim\text{Pois}(h)$ can be 1700, 800, 1200, where Pois is the Poisson distribution. The rationale is that larger numbers do not result in more difficult problems, but more scientific numbers do. An example would be comparing the numbers 150,000 and 172,568.
@@
The more restrictive a method is, the closer the difficulty levels and coherence are between the an adversarial example and its original version. Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with all constraints to maintain the original difficulty and coherence. Although generations from $M1$ and $M2$ may not fit in the educational context, we are interested in studying LLMs’ math-solving abilities by simply changing the numeric values. We discuss how different methods impact model performance in §4.2. We describe each generation method in more detail in Appendix D and present the generated examples based on three methods in Table 5.
@@ 4.1 Experimental Setup
We generate problem variants from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015). Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced from math worksheets used by elementary school students to practice math problems (Roy et al., 2015). [^4] While MWPs exhibit different levels of difficulty, we focus on MWPs at this particular difficulty level as a proof of concept in this work.
Models
We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama 2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4 (OpenAI, 2023) and GPT-3.5 (OpenAI, 2022). [^5] Specifically, we use GPT-4-0125-preview and GPT-3.5-turbo-0125. For simplicity, we refer to them as GPT-4 and GPT-3.5 for the rest of the paper. The model selection was largely based on the LLMs’ math-solving ability. Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama 2 and Mistral, and the API-based GPT models.
Metrics
We follow the previous adversarial attack literature to measure the attacks, which looks for at least one perturbation that fools the model (Croce et al., 2020; Pruthi et al., 2019; Jia & Liang, 2017). A problem is considered incorrect if it has at least one incorrect variation. Such variation is also known as an adversarial example. Given a set of original problem-answer pairs $P$ and a LLM $L$, we define: [^6] OA, AA, ASR are reported in % in this work.
  • Original Accuracy (OA): the accuracy of $L$ on the original problems,
    $\text{OA}(L)=\frac{\sum_{(x,y)\in P}\mathbf{1}\{L(x)=y\}}{|P|}.$ (3)
  • Attack Accuracy (AA): the accuracy of $L$ on the adversarial modifications,
    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right]}{|P|}.$ (4)
    $\text{ASR}(L)=\frac{\text{OA}(L)-\text{AA}(L)}{\text{OA}(L)}.$ (5)
@@
Mistral 7B 37 0 100 0 100 0 100 29 0 100 0 100 0 100
MetaMath 7B 100 74 26 10 90 0 100 95 28 71 9 91 0 100
WizardMath 13B 89 20 78 5 94 0 100 89 11 88 2 98 0 100
Vicuna 13B 76 4 95 1 99 0 100 60 0 100 0 100 0 100
Llama 2 13B 12 0 100 0 100 0 100 10 0 100 0 100 0 100
CodeLlama 34B 11 0 100 0 100 0 100 6 0 100 0 100 0 100
MetaMath 70B 99 86 13 30 70 0 100 98 50 49 17 83 0 100
GPT-3.5 97 74 24 47 52 0 100 91 52 43 31 66 0 100
Average 65.1 32.3 66.9 11.6 88.1 0 100 59.8 17.6 81.4 7.4 92.3 0 100
Table 1: Main Attacks We generate MWPs based on different generation methods. Overall, simply changing the numeric values in the problems causes a consistent performance drop across all LLMs, even with the most restrictive generation method.
@@
As a baseline, we examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs. Zhou et al. (2023) freeze the logical entities and iteratively swap each unfrozen token with similar tokens that cause LLMs to fail. They then validate the adversarial example with human manual checking. Zhou et al. (2023) created the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith with their 300 rephrased versions. We report the results in Appendix E, observing ASR for 4 out of the 7 models; however, some models even show improved accuracy given these adversarial examples, suggesting that such attacks may not generalize well across different LLMs.

### v3

# Adversarial Math Word Problem Generation

@@ Abstract
Large language models (LLMs) have significantly transformed the educational landscape. As current plagiarism detection tools struggle to keep pace with LLMs’ rapid advancements, the educational community faces the challenge of assessing students’ true problem-solving abilities in the presence of LLMs. In this work, we explore a new paradigm for ensuring fair evaluation—generating adversarial examples which preserve the structure and difficulty of the original questions aimed for assessment, but are unsolvable by LLMs. Focusing on the domain of math word problems, we leverage abstract syntax trees to structurally generate adversarial examples that cause LLMs to produce incorrect answers by simply editing the numeric values in the problems. We conduct experiments on various open- and closed-source LLMs, quantitatively and qualitatively demonstrating that our method significantly degrades their math problem-solving ability. We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models. Additionally, we conduct automatic analysis to investigate the cause of failure, providing further insights into the limitations of LLMs.  [^1] Data and code are available: https://github.com/ruoyuxie/adversarial_mwps_generation
Adversarial Math Word Problem Generation
Roy Xie      Chengxuan Huang      Junlin Wang      Bhuwan Dhingra Duke University {ruoyu.xie, jonathan.huang, junlin.wang2}@duke.edu {bdhingra}@cs.duke.edu
@@ 1 Introduction
Figure 1: Method Overview: Given a MWP that an LLM can correctly solve, our method first transforms it into Python code. The Python code then is converted into an AST representation, which is used to generate adversarial problems by modifying the numeric values in a controllable manner. We place constraints on the nodes of the AST to ensure that the modified problem maintains the same difficulty level as the original problem. Despite this, we find that the resulting adversarial examples cause LLMs to predict incorrect answers.
@@
In this work, we evaluate several LLMs and demonstrate the effectiveness of our method by achieving a significant attack success rate (ASR). Our approach outperforms the previous rephrasing attack by an average of 62% ASR. We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models (e.g., GPT-4) by reducing API request calls by 90% while achieving high performance. We conduct human evaluations to verify that our generated problems indeed preserve the original coherence and difficulty. Furthermore, we perform a regression analysis and find that our adversarial examples exploit different weaknesses of each model, offering valuable insights into LLM’s limitation.
@@ 2 Background and Related Work
Our work also closely relates to LLM’s math reasoning ability (Yu et al., 2023; Xu et al., 2023). Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023a), such as chain-of-thought (CoT) (Wei et al., 2022). Converting the MWPs’ solving steps into symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023). [^2] While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve. Therefore, we try to minimize prompt engineering by using a unified prompt with zero-shot CoT for all evaluations. The prompt can be found in the appendix A.2.
@@
Adversarial attacks on LLMs involve modifying prompts to alter their behavior (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023). In this work, we modify MWPs to cause LLMs to output incorrect answers. Bubeck et al. (2023) conducted limited memorization tests on MWPs by randomly changing numeric values, suggesting that state-of-the-art LLMs do not solely rely on memorization but apply general solution methods. However, their study’s sample size was relatively small, and we demonstrate that modifying numbers in MWPs causes LLMs to fail on a larger scale. On the other hand, Zhou et al. (2023) rephrased MWPs by changing words while preserving numeric values. This approach risks altering the original context and introducing inconsistencies and problematic content (see Appendix B), requiring extensive human validation. Therefore, we focus on altering the numerical values in this work and preserve the underlying logic.
@@ 3 Methodology
MWPs are presented in natural language, which creates challenges for systematic structural and syntactic modifications. In this section, we describe our approach to generating adversarial MWPs. An overview of our method can be found in Figure 1. We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer. We define $G$ as a ground-truth function that computes the correct answer for a math problem and $F$ as a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise. [^3] In our case $G$ is generated Python code. We denote the set of all possible adversarial modifications to it as:
@@ 3.1 Mapping From MWPs to Code to Tree
To structurally modify MWPs, we first use GPT-4 to generate the Python code that reflects the solution steps given the problem and its final answer. [^4] This is an one-time operation. We only require the problem and its final answer, which is almost always given in the context of math problems. The prompt is in the Appendix A.1. Next, we utilize the AST, a tree structure for programming language code, to convert the generated Python code into a tree representation for controllable new problem generation. We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement. The final print statement that outputs the answer is the root of the tree. Each AST has mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes. We discuss the nodes in ASTs in detail in Appendix C and conduct analysis on the quality of generated Python code and ASTs in §5.
@@ 3.2 Adversarial Example Generation
In an adversarial setting, attackers can perform multiple rounds of attack on the system, and any successful attack signifies the possible flaws of the system (Carlini and Wagner, 2017). Similarly, we generate adversarial examples which are different versions of the original MWPs to attack LLMs.
@@
It’s essential to maintain the original difficulty and coherence of the MWPs despite changes in their numeric values. For instance, changing one-digit multiplication to four-digit multiplication significantly alters the problem’s complexity. To maintain mathematical logic intact, the order of magnitude of numbers in the problem should also remain unchanged. For example, changing a problem from “Jack ate 2 out of a total of 5 apples” to “Jack ate 5 out of a total of 2 apples” not only introduces the concept of negative numbers but also alters the problem’s logical structure and coherency. Such modifications can create unrealistic scenarios, potentially confusing students and leading to ineffective learning experiences Vilenius-Tuohimaa et al. (2008).
@@
To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node in the AST to control the generation quality. Given the original node value $h$ and its new value $h^{\prime}$, a newly generated problem is valid if and only if $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes. A list of node constraints is:
@@
The node constraints ensure the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated. Different values for the variable nodes can also lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$. To mitigate this, we propose three generation methods to distinguish generated adversarial examples with different levels of difficulty. We decompose the values $h\approx a\times 10^{b}$, where $a$ is an integer that is not divisible by 10 and with at most $c$ digits. Depending on the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
  • $M1$ Free Generation: Allowing the generated problem to have a wide range of numbers with minimal constraints, with each $h^{\prime}$ to be $a^{\prime}\times 10^{b}$. Regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$. If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$.
  • $M2$ Count of Digits: Constraining the generated value $h^{\prime}$ has the same number of digits as $h$. For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264. This ensures $h^{\prime}$ is always within a reasonable range relative to $h$ and maintains similar problem difficulty levels. For more variability, $h^{\prime}$ can range from 1 to 99 if $h$ has only one digit. For decimal variable values like $h=1.25$, we use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
  • $M3$ Count of Scientific Numbers: Constraining the new value shares a similar scientific digit count and is within a similar range to the original value. For example, for $h=1500$, $h^{\prime}\sim\text{Pois}(h)$ can be 1700, 800, 1200, where Pois is the Poisson distribution. The rationale is that larger numbers do not result in more difficult problems, but more scientific numbers do. An example would be comparing the numbers 150,000 and 172,568.
@@
The more restrictive a method is, the closer the difficulty levels and coherence remain between an adversarial example and its original version. Thus, we use $M3$ as our main generation method since it ensures that adversarial examples adhere to all constraints, maintaining original difficulty and coherence. While $M1$ and $M2$ generations may not fit the educational context, we aim to study LLMs’ math-solving abilities by simply altering numeric values. We discuss how different methods impact model performance in §4.2. We present detailed descriptions of each generation method in Appendix D and their generated examples in Table 6.
@@ 4.1 Experimental Setup
We generate problem variants from GSM8K (Cobbe et al., 2021) and MultiArith (Roy and Roth, 2015). Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced directly from math worksheets used by elementary school students to practice math problems (Roy et al., 2015). [^5] While MWPs exhibit different levels of difficulty, we focus on MWPs at this particular difficulty level as a proof of concept in this work.
@@
Mistral 7B 37.0 0.0 100.0 0.0 100.0 0.0 100.0 29.0 0.0 100.0 0.0 100.0 0.0 100.0
MetaMath 7B 100.0 74.0 26.0 10.0 90.0 0.0 100.0 95.0 28.0 71.0 9.0 91.0 0.0 100.0
Llama 3 8B 17.0 0.0 100.0 0.0 100.0 0.0 100.0 21.0 0.0 100.0 0.0 100.0 0.0 100.0
Llama 2 13B 12.0 0.0 100.0 0.0 100.0 0.0 100.0 10.0 0.0 100.0 0.0 100.0 0.0 100.0
WizardMath 13B 89.0 20.0 78.0 5.0 94.0 0.0 100.0 89.0 11.0 88.0 2.0 98.0 0.0 100.0
Vicuna 13B 76.0 4.0 95.0 1.0 99.0 0.0 100.0 60.0 0.0 100.0 0.0 100.0 0.0 100.0
CodeLlama 34B 11.0 0.0 100.0 0.0 100.0 0.0 100.0 6.0 0.0 100.0 0.0 100.0 0.0 100.0
MetaMath 70B 99.0 86.0 13.0 30.0 70.0 0.0 100.0 98.0 50.0 49.0 17.0 83.0 0.0 100.0
GPT-3.5 97.0 74.0 24.0 47.0 52.0 0.0 100.0 91.0 52.0 43.0 31.0 66.0 0.0 100.0
Average 60.0 28.7 70.7 10.3 78.3 0.0 100.0 55.4 15.9 83.4 6.6 93.1 0.0 100.0
Table 1: Main Attacks: Performance of three different generation methods. Simply changing numeric values consistently cause performance drop across all LLMs, even with the most restrictive generation method.
Models
We conduct experiments on the following open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama 3 8B (AI@Meta, 2024), Llama 2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), and CodeLlama (Roziere et al., 2023). Additional, we evaluate two closed-source models: GPT-4 (OpenAI, 2023) and GPT-3.5 (OpenAI, 2022). [^6] GPT-4-0125-preview and GPT-3.5-turbo-0125. The model selection was largely based on the LLMs’ math-solving ability. Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama 3 and Mistral, and the API-based GPT models.
Metrics
We follow the previous adversarial attack literature to measure the attacks, which looks for at least one perturbation that fools the model (Croce et al., 2020; Pruthi et al., 2019; Jia and Liang, 2017). A problem is considered incorrect if it has at least one incorrect variation, also known as an adversarial example. Given a set of original problem-answer pairs $P$ and a LLM $L$, we define: [^7] OA, AA, ASR are reported in % in this work.
  • Original Accuracy (OA): the accuracy of $L$ on the original problems,
    $\text{OA}(L)=\frac{\sum_{(x,y)\in P}\mathrm{1}\{L(x)=y\}}{|P|}.$ (3)
  • Attack Accuracy (AA): given an indicator function $I_{xy}$:
    $I_{xy}=\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right],$ (4)
    $L$’s accuracy on the adversarial examples:
    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}I_{xy}}{|P|}.$ (5)
    $\text{ASR}(L)=\frac{\text{OA}(L)-\text{AA}(L)}{\text{OA}(L)}.$ (6)
@@
As a baseline, we evaluate a common rephrasing approach focusing on modifying words within MWPs. Zhou et al. (2023) freeze the logical entities and iteratively swap each unfrozen token with similar tokens that cause LLMs to fail. They then validate the adversarial example with human manual checking. Zhou et al. (2023) created the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith with their 300 rephrased versions. We report the results in Appendix E, observing ASR for 4 out of the 7 models; however, some models even show improved accuracy given these adversarial examples, suggesting that such attacks may not generalize well across different LLMs.